OpenStudy (vortish):

the perimeter of a rectangle is 38 ft the length is 3 ft longer than the width. write a system of linear equations and solve the resulting system let x be the length and y b the width 2x +2y = ? x= y + ?

OpenStudy (campbell_st):

2x + 2y = 38 then x = 19 - y

OpenStudy (vortish):

campbell_st can you expand on that answer please

OpenStudy (campbell_st):

well you know that x = 3 + y that is the rectangle is 3 units longer than the width then the equations are x = 3 + y (1) x = 19 - y (2) equate the equations 3 + y = 19 - y add y to both sides 3 + 2y = 19 subtract 3 from both sides 2y = 16 or y = 8 then the width is 8 units so length = 11 units found by substituting into equation 2 above